OpenStudy (anonymous):

-6x7=-5 and the answer is 2. Need to show the work help me.

OpenStudy (anonymous):

is that (-6) times (7) or (-6x) (7)

OpenStudy (anonymous):

No it's -6x like the letter +7=-5

OpenStudy (anonymous):

ok -x6+7 = -5 subract 7 from both sides -6x = -12 divide both sides by -6 -6x/-6 = -12/-6 x = 2

OpenStudy (anonymous):

cool thank u soo much I'm going to bed now, but I'll join u guys in the morning or in the afternoon when I get home from school. I'm also doing diagrams and such.

OpenStudy (anonymous):

cool thank u soo much I'm going to bed now, but I'll join u guys in the morning or in the afternoon when I get home from school. I'm also doing diagrams and such.